OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 Solve for x: −2(x + 3) = −2(x + 1) − 4 One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. *x = 0.6x + 384* How many people who downloaded songs from the site that month downloaded Q Sam's latest single?

OpenStudy (adilalvi):

3x+3=-2x+1+6 3x+2x=7-3 5x=4 x=5/4 answer (b); -2x-6=-2x-2-4 -2x+2x=-6+6 0=0 answer

OpenStudy (adilalvi):

sorry 1st answer is x=4/5 not 5/4

OpenStudy (anonymous):

solve for x: 3(x + 1) = −2(x − 1) + 6 3x+3=-2x+2+6 3x+2x=-3+2+6 5x=-1+6 x=1

OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2-4 -2x+2x=6-2-4 x=4-4 x=0